OpenStudy (anonymous):

how do you convert y=9+x to standard form

OpenStudy (ksaimouli):

what do u mean by S F

OpenStudy (ksaimouli):

m=1 yint= 9

OpenStudy (anonymous):

I need it to go to his form ::: Ax+By=C

OpenStudy (ksaimouli):

c=9 A=1 B=1

OpenStudy (ksaimouli):

0r c=18 A=2 B=2

OpenStudy (zzr0ck3r):

\(y=mx+b\) you are in this form already (addition is commutative) you have \(y=9+x = x+9 = 1*x+9\)